OpenStudy (anonymous):

What is the approximate area of a circle with a radius of 28 centimeters? Use π ≈ 3.14. A. 2461.8 square centimeters B. 175.8 square centimeters C. 4923.5 square centimeters D. 87.9 square centimeters

OpenStudy (dejmccray):

C=pi*2*R=3.14*2*28= 2*28= 56*3.14=175.84 which is b